OpenStudy (anonymous):

A cube has a side length of 4 m. What is the volume of the cube? I THINK IT IS 64, DOES ANYONE ELSE AGREE???

OpenStudy (anonymous):

@Directrix , i think it is 64, what do u say??

OpenStudy (anonymous):

V = 4(3) = 64

OpenStudy (anonymous):

the three is small

OpenStudy (anonymous):

It is 64. The volume of a cube is its width times its length times its height. Since those are all 4, 4 x 4 x 4 = 64.

OpenStudy (anonymous):

I think its 64

OpenStudy (gabylovesyou):

Volume of a cube is s^3 so... if you replace s with 4 you get 4^3 4 * 4 * 4 = ??

OpenStudy (anonymous):

I agree because volume equals l * W times H and 4 * 4 * 4 = 64

OpenStudy (anonymous):

in that case, thx guys, it is 64

OpenStudy (gabylovesyou):

:) good job ..

Directrix (directrix):

You want to find the cube of 4. 4^3 = 4*4*4 = ?

OpenStudy (anonymous):

yep so you answered ur own question good job @upandup1445

OpenStudy (anonymous):

good job you got it